# A Prompt-Engineered Large Language Model, Deep Learning Workflow for Materials Classification

### v1

@@ Discussion
We emphasize that classification cannot be successfully achieved using direct questioning alone (as seen in Supplementary Figs. S8-S10). Therefore, as part of our workflow, we designed a BERT-based classification model to categorize materials using text generated by the LLM. In practice, employing a classification model is a necessary step at this juncture, as directly querying the LLM yielded a 0 accuracy rate.
The general workflow here offers a novel approach to address a significant challenge in “AI for material science”, where datasets are often small. By utilizing a pre-trained BERT model, such as MatSciBERT optimized for material science, we can fine-tune it (pre-trained model) with textual data from prompt-engineered LLM to achieve high accuracy in classification problem (up to 99.0% for the entire dataset).
We propose a general approach for materials classification and demonstrate its efficiency in an example case of the classification problem in MGs. Our general workflow can be applied to any type of material applications, provided that labeled material samples are available. The findings point towards a future where the integration of advanced language models and domain-specific fine-tuning could revolutionize the material classification process, especially in cases with sparse data. We envision that LLMs may eventually serve as “world models” [49], encapsulating knowledge across diverse materials and domains. Consequently, such models based on text for classification, prediction, or even generating new materials will become potential tools for addressing challenges in material design. A sufficiently robust LLM could potentially provide desired classification outcomes directly through appropriate prompting and significantly streamline the process of materials discovery and design.
@@ Methods
In processing the dataset for metallic glasses, we curated our initial compilation of 8,415 entries by excluding alloys comprising multiple categories. This refinement resulted in a dataset encompassing 5,577 single-class alloys, along with their respective classifications. The aim was to construct an unambiguous dataset, thereby facilitating the accuracy of analysis and model training.
To implement prompt engineering, we harnessed the capabilities of the langchain library. We designed prompts with varying templates in advance, establishing a “Composition” variable to dynamically cycle through the list of alloy compositions. This approach enabled us to systematically generate textual data in a large scale, ensuring consistency and efficiency in our data preparation workflow.
For the ingestion of data into the MgBERT model, we divided the 5,577 data points into a training and test set at a ratio of 80:20. This arrangement yielded 4,460 entries for training constituting 80% of the dataset, and 1,117 entries for testing purposes. It was imperative that the MgBERT model was trained comprehensively, indicating that the data allocation is reasonable.
Regarding the evaluation of datasets, while the MgBERT model was assessed using the designated test set, the machine learning models underwent evaluation using a different approach because they can not receive textual data. We employed the Matminer and Pymatgen tools to convert the chemical formulas of alloys into “composition features” as defined within Matminer’s framework. However, it is noteworthy that three entries failed to transform correctly. Consequently, the test set for evaluation purposes comprised of 1,116 entries, reducing the overall dataset to 5,574 records. Compared to MgBERT, there were three fewer data sets, which had little impact on accuracy comparison.
Large language model and prompt engineering
@@
In this work, we used the Google Gemini-pro large language model, which is currently free and surpasses human experts in massive multi-task language understanding [50]. As discussed earlier, our research harnesses the automated orchestration capabilities of the langchain library to script the bulk generation of prompts and textual data. This scripting forms an integral part of a generic workflow and offers a high degree of adaptability. The prompt templates we have developed are not rigid constructs; rather, they are designed to be inherently flexible, allowing for arbitrary modifications tailored to the specific demands and objectives of researchers. This level of customization is critical in the context of materials science where the nuances of the subject matter can vary significantly from one study to another. The ability to fine-tune prompts to align with particular research questions or datasets ensures that the language model can be effectively leveraged to generate effective and contextually relevant textual data outputs.
@@
Our investigation sought to demonstrate the effectiveness of prompt engineering by creating a range of alternative prompt templates to act as comparative benchmarks. These templates comprised one that integrated a chain-of-thoughts (CoT) approach with a few-shot learning paradigm (denoted as “few-shot with CoT”), another that depended exclusively on few-shot learning examples (denoted as “few-shot”), a third that was based on zero-shot learning scenarios (denoted as “direct inquiry”), and our custom-designed “MetalPrompt” from our established workflow. By leveraging these diverse templates, we crafted specific prompts that were subsequently employed to interrogate the Gemini-pro model. The underlying rationale for this varied prompt strategy was to gather and juxtapose the accuracy metrics derived from different prompting methodologies.
@@
Here we used MatSciBERT as the basic model and implemented a subsequent classifier function with MgBERT embedding to build MgBERT. Accuracy and cross entropy are used to evaluate the effectiveness of the model. For the classification principle of MgBERT, we used the value of [CLS] token in the last layer of the attention section of MgBERT as the output to aggregate the information of the entire model. Then, through a classifier, we compress the value of the token into a vector of length 3, and label 0 as BMG, 1 as Ribbon, and 2 as Non-ribbon. Then we calculate the maximum position of the output vector value and locate it to the final classification result. The overall expenses associated with developing MgBERT comprise $46 USD for training on Nvidia V100 (30 hours and $1.53 USD per hour) and $1.5 USD for 1.89 million input and 2.14 million output tokens (according to the official pricing of Gemini-pro on January 29, 2024). Upon applying MgBERT for inference on 1,000 different compositions, the cost is as low as $0.3 USD. This cost-effective nature of training and implementing MgBERT enables researchers to adopt this general approach across various materials.
@@
To elucidate the relationship between input and output, we used the LIME (local interpretable model agnostic explanations) technique, which is widely used to explain the predictions of machine learning models by constructing the relationship function via interpolation sampling. Here we displayed the top 25 important features and performed 1,000 interpolation samples for function fitting. Although our parameters were initially set to select the top 50 features, due to visualization constraints, we chose the top 25 features for display.
For the explanation at the attention level, we used the bertviz library to complete the attention score flow diagram for 12 layers and 12 attention heads. To attribute the importance of the [CLS] token in the final attention layer, we first averaged the attention score of the 12 attention heads in the last layer. Then, we divided the input text words into three layers based on their positions, namely, composition, element, and alloy layer. To evaluate the importance of each layer, we calculated the sum of attention scores within that layer, weighted by the token length ratio, and then divided by the total token length.
@@
The method section provides the models and algorithms employed in this study, while specific parameter implementations can be found in the supplementary notes in supplementary information. The data used in this work is available in the supplementary data, including the metallic glasses dataset, alloy description files generated by large language model Gemini-pro, MgBERT classification model training logs, and data used for figure plotting.
@@
All the codes used in the paper will be made available on GitHub upon the acceptance of manuscript, and these codes can also be acquired upon reasonable requests.
ReferencesReferences 1-50 are for the main text.
- Louie et al. [2021] S. G. Louie, Y.-H. Chan, F. H. da Jornada, Z. Li, and D. Y. Qiu, Discovering and understanding materials through computation, Nature Materials 20, 728 (2021).
- Merchant et al. [2023] A. Merchant, S. Batzner, S. S. Schoenholz, M. Aykol, G. Cheon, and E. D. Cubuk, Scaling deep learning for materials discovery, Nature 624, 80 (2023).
- Li et al. [2020] J. Li, K. Lim, H. Yang, Z. Ren, S. Raghavan, P.-Y. Chen, T. Buonassisi, and X. Wang, Ai applications through the whole life cycle of material discovery, Matter 3, 393 (2020).
- Raabe et al. [2023] D. Raabe, J. R. Mianroodi, and J. Neugebauer, Accelerating the design of compositionally complex materials via physics-informed artificial intelligence, Nature Computational Science 3, 873 (2023).
- Nippa et al. [2023] D. F. Nippa, K. Atz, R. Hohler, A. T. Müller, A. Marx, C. Bartelmus, G. Wuitschik, I. Marzuoli, V. Jost, J. Wolfard, et al., Enabling late-stage drug diversification by high-throughput experimentation with geometric deep learning, Nature Chemistry , 1 (2023).
- Tropsha et al. [2023] A. Tropsha, O. Isayev, A. Varnek, G. Schneider, and A. Cherkasov, Integrating qsar modelling and deep learning in drug discovery: the emergence of deep qsar, Nature Reviews Drug Discovery , 1 (2023).
- Weiss et al. [2023] T. Weiss, E. Mayo Yanes, S. Chakraborty, L. Cosmo, A. M. Bronstein, and R. Gershoni-Poranne, Guided diffusion for inverse molecular design, Nature Computational Science 3, 873 (2023).
- Noé et al. [2020] F. Noé, A. Tkatchenko, K.-R. Müller, and C. Clementi, Machine learning for molecular simulation, Annual review of physical chemistry 71, 361 (2020).
- Wigh et al. [2022] D. S. Wigh, J. M. Goodman, and A. A. Lapkin, A review of molecular representation in the age of machine learning, Wiley Interdisciplinary Reviews: Computational Molecular Science 12, e1603 (2022).
- Choudhary et al. [2022] K. Choudhary, B. DeCost, C. Chen, A. Jain, F. Tavazza, R. Cohn, C. W. Park, A. Choudhary, A. Agrawal, S. J. Billinge, et al., Recent advances and applications of deep learning methods in materials science, npj Computational Materials 8, 59 (2022).
- Antoniuk et al. [2023] E. R. Antoniuk, G. Cheon, G. Wang, D. Bernstein, W. Cai, and E. J. Reed, Predicting the synthesizability of crystalline inorganic materials from the data of known material compositions, npj Computational Materials 9, 155 (2023).
- Noh et al. [2020] J. Noh, G. H. Gu, S. Kim, and Y. Jung, Machine-enabled inverse design of inorganic solid materials: promises and challenges, Chemical Science 11, 4871 (2020).
- Xiao et al. [2023] H. Xiao, R. Li, X. Shi, Y. Chen, L. Zhu, X. Chen, and L. Wang, An invertible, invariant crystal representation for inverse design of solid-state materials using generative deep learning, Nature Communications 14, 7027 (2023).
- Steinberger et al. [2019] D. Steinberger, H. Song, and S. Sandfeld, Machine learning-based classification of dislocation microstructures, Frontiers in Materials 6, 141 (2019).
- Jain et al. [2013] A. Jain, S. P. Ong, G. Hautier, W. Chen, W. D. Richards, S. Dacek, S. Cholia, D. Gunter, D. Skinner, G. Ceder, and K. A. Persson, Commentary: The Materials Project: A materials genome approach to accelerating materials innovation, APL Materials 1, 011002 (2013).
- Zagorac et al. [2019] D. Zagorac, H. Müller, S. Ruehl, J. Zagorac, and S. Rehme, Recent developments in the Inorganic Crystal Structure Database: theoretical crystal structure data and related features, Journal of Applied Crystallography 52, 918 (2019).
- Curtarolo et al. [2012] S. Curtarolo, W. Setyawan, G. L. Hart, M. Jahnatek, R. V. Chepulskii, R. H. Taylor, S. Wang, J. Xue, K. Yang, O. Levy, M. J. Mehl, H. T. Stokes, D. O. Demchenko, and D. Morgan, Aflow: An automatic framework for high-throughput materials discovery, Computational Materials Science 58, 218 (2012).
- Vu et al. [2023] T.-S. Vu, M.-Q. Ha, D.-N. Nguyen, V.-C. Nguyen, Y. Abe, T. Tran, H. Tran, H. Kino, T. Miyake, K. Tsuda, et al., Towards understanding structure–property relations in materials with interpretable deep learning, npj Computational Materials 9, 215 (2023).
- Thirunavukarasu et al. [2023] A. J. Thirunavukarasu, D. S. J. Ting, K. Elangovan, L. Gutierrez, T. F. Tan, and D. S. W. Ting, Large language models in medicine, Nature medicine 29, 1930 (2023).
- Kasneci et al. [2023] E. Kasneci, K. Seßler, S. Küchemann, M. Bannert, D. Dementieva, F. Fischer, U. Gasser, G. Groh, S. Günnemann, E. Hüllermeier, et al., Chatgpt for good? on opportunities and challenges of large language models for education, Learning and individual differences 103, 102274 (2023).
- Romera-Paredes et al. [2023] B. Romera-Paredes, M. Barekatain, A. Novikov, M. Balog, M. P. Kumar, E. Dupont, F. J. Ruiz, J. S. Ellenberg, P. Wang, O. Fawzi, et al., Mathematical discoveries from program search with large language models, Nature 625, 468 (2023).
- Pereira et al. [2023] J. Pereira, R. Fidalgo, R. Lotufo, and R. Nogueira, Visconde: Multi-document qa with gpt-3 and neural reranking, in European Conference on Information Retrieval (Springer, 2023) pp. 534–543.
- Chen et al. [2024] D. Chen, S. Song, Q. Yu, Z. Li, W. Wang, F. Xiong, and B. Tang, Grimoire is all you need for enhancing large language models (2024), arXiv:2401.03385 .
- Xie et al. [2023] T. Xie, Y. Wa, W. Huang, Y. Zhou, Y. Liu, Q. Linghu, S. Wang, C. Kit, C. Grazian, and B. Hoex, Large language models as master key: Unlocking the secrets of materials science with gpt (2023), arXiv:2304.02213 .
- Zheng et al. [2023] Z. Zheng, O. Zhang, C. Borgs, J. T. Chayes, and O. M. Yaghi, Chatgpt chemistry assistant for text mining and the prediction of mof synthesis, Journal of the American Chemical Society 145, 18048 (2023).
- Min et al. [2023] B. Min, H. Ross, E. Sulem, A. P. B. Veyseh, T. H. Nguyen, O. Sainz, E. Agirre, I. Heintz, and D. Roth, Recent advances in natural language processing via large pre-trained language models: A survey, ACM Computing Surveys 56, 1 (2023).
- Devlin et al. [2018] J. Devlin, M.-W. Chang, K. Lee, and K. Toutanova, Bert: Pre-training of deep bidirectional transformers for language understanding (2018), arXiv:1810.04805 .
- Morgan and Jacobs [2020] D. Morgan and R. Jacobs, Opportunities and challenges for machine learning in materials science, Annual Review of Materials Research 50, 71 (2020).
- Deng et al. [2009] J. Deng, W. Dong, R. Socher, L.-J. Li, K. Li, and L. Fei-Fei, Imagenet: A large-scale hierarchical image database, in 2009 IEEE conference on computer vision and pattern recognition (Ieee, 2009) pp. 248–255.
- Fan et al. [2023] X. Fan, S. Chen, B. Steingrimsson, Q. Xiong, W. Li, and P. K. Liaw, Dataset for fracture and impact toughness of high-entropy alloys, Scientific Data 10, 37 (2023).
- Zhang et al. [2023a] Z. Zhang, H. Tang, and Z. Xu, Fatigue database of complex metallic alloys, Scientific Data 10, 447 (2023a).
- Gorsse et al. [2018] S. Gorsse, M. Nguyen, O. N. Senkov, and D. B. Miracle, Database on the mechanical properties of high entropy alloys and complex concentrated alloys, Data in brief 21, 2664 (2018).
- Li et al. [2023] X. Li, X. Hu, X. Chen, J. Fan, Z. Zhao, J. Wu, H. Wang, and Q. Dai, Spatial redundancy transformer for self-supervised fluorescence image denoising, Nature Computational Science 3, 1067 (2023).
- Zhou et al. [2023] G. Zhou, Z. Gao, Q. Ding, H. Zheng, H. Xu, Z. Wei, L. Zhang, and G. Ke, Uni-mol: a universal 3d molecular representation learning framework (2023).
- Varshneya and Mauro [2019] A. K. Varshneya and J. C. Mauro, Chapter 1 - introduction, in Fundamentals of Inorganic Glasses (Third Edition), edited by A. K. Varshneya and J. C. Mauro (Elsevier, 2019) third edition ed., pp. 1–18.
- Halim et al. [2021] Q. Halim, N. A. N. Mohamed, M. R. M. Rejab, W. N. W. A. Naim, and Q. Ma, Metallic glass properties, processing method and development perspective: a review, The International Journal of Advanced Manufacturing Technology 112, 1231 (2021).
- Liu et al. [2023a] G. Liu, S. Sohn, C. S. O’Hern, A. C. Gilbert, and J. Schroers, Effective subgrouping enhances machine learning prediction in complex materials science phenomena: Inoue’s subgrouping in discovering bulk metallic glasses, Acta Materialia 265, 119590 (2023a).
- Li et al. [2017] Y. Li, S. Zhao, Y. Liu, P. Gong, and J. Schroers, How many bulk metallic glasses are there?, ACS combinatorial science 19, 687 (2017).
- Zhou et al. [2022] Z. Zhou, Y. Shang, and Y. Yang, A critical review of the machine learning guided design of metallic glasses for superior glass-forming ability, Journal of Materials Informatics 2, 1 (2022).
- Ward et al. [2018] L. Ward, S. C. O’Keeffe, J. Stevick, G. R. Jelbert, M. Aykol, and C. Wolverton, A machine learning approach for engineering bulk metallic glass alloys, Acta Materialia 159, 102 (2018).
- Rodriguez et al. [2023] A. D. Rodriguez, K. R. Dearstyne, and J. Cleland-Huang, Prompts matter: Insights and strategies for prompt engineering in automated software traceability, in 2023 IEEE 31st International Requirements Engineering Conference Workshops (REW) (IEEE, 2023) pp. 455–464.
- Gupta et al. [2022] T. Gupta, M. Zaki, N. A. Krishnan, and Mausam, Matscibert: A materials domain language model for text mining and information extraction, npj Computational Materials 8, 102 (2022).
- Chakraborty et al. [2017] S. Chakraborty, R. Tomsett, R. Raghavendra, D. Harborne, M. Alzantot, F. Cerutti, M. Srivastava, A. Preece, S. Julier, R. M. Rao, et al., Interpretability of deep learning models: A survey of results, in 2017 IEEE smartworld, ubiquitous intelligence & computing, advanced & trusted computed, scalable computing & communications, cloud & big data computing, Internet of people and smart city innovation (IEEE, 2017) pp. 1–6.
- Ribeiro et al. [2016] M. T. Ribeiro, S. Singh, and C. Guestrin, ”Why Should I Trust You?”: Explaining the Predictions of Any Classifier (2016), arXiv:1602.04938 .
- Inoue [2000] A. Inoue, Stabilization of metallic supercooled liquid and bulk amorphous alloys, Acta materialia 48, 279 (2000).
- Ren et al. [2018] F. Ren, L. Ward, T. Williams, K. J. Laws, C. Wolverton, J. Hattrick-Simpers, and A. Mehta, Accelerated discovery of metallic glasses through iteration of machine learning and high-throughput experiments, Science advances 4, eaaq1566 (2018).
- Johnson et al. [2016] W. Johnson, J. Na, and M. Demetriou, Quantifying the origin of metallic glass formation, Nature communications 7, 10313 (2016).
- Wang [2009] W. Wang, Bulk metallic glasses with functional physical properties, Advanced Materials 21, 4524 (2009).
- Dawid and LeCun [2023] A. Dawid and Y. LeCun, Introduction to latent variable energy-based models: A path towards autonomous machine intelligence (2023), arXiv:2306.02572 .
- Hendrycks et al. [2021] D. Hendrycks, C. Burns, S. Basart, A. Zou, M. Mazeika, D. Song, and J. Steinhardt, Measuring massive multitask language understanding (2021), arXiv:2009.03300 .
- Joiner [2018] I. A. Joiner, Chapter 1 - artificial intelligence: Ai is nearby, in Emerging Library Technologies, Chandos Information Professional Series, edited by I. A. Joiner (Chandos Publishing, 2018) pp. 1–22.
- Niu et al. [2021] Z. Niu, G. Zhong, and H. Yu, A review on the attention mechanism of deep learning, Neurocomputing 452, 48 (2021).
- Wei et al. [2023] J. Wei, X. Wang, D. Schuurmans, M. Bosma, B. Ichter, F. Xia, E. Chi, Q. Le, and D. Zhou, Chain-of-thought prompting elicits reasoning in large language models (2023), arXiv:2201.11903 .
- Ray [2023] P. P. Ray, Chatgpt: A comprehensive review on background, applications, key challenges, bias, ethics, limitations and future scope, Internet of Things and Cyber-Physical Systems 3, 121 (2023).
- Helm et al. [2020] J. M. Helm, A. M. Swiergosz, H. S. Haeberle, J. M. Karnuta, J. L. Schaffer, V. E. Krebs, A. I. Spitzer, and P. N. Ramkumar, Machine learning and artificial intelligence: definitions, applications, and future directions, Current reviews in musculoskeletal medicine 13, 69 (2020).
- Almeida and Xexeo [2023] F. Almeida and G. Xexeo, Word embeddings: A survey (2023), arXiv:1901.09069 .
- Zheng and Casari [2018] A. Zheng and A. Casari, Feature Engineering for Machine Learning: Principles and Techniques for Data Scientists, 1st ed. (O’Reilly Media, Inc., 2018).
- Svozil et al. [1997] D. Svozil, V. Kvasnicka, and J. Pospichal, Introduction to multi-layer feed-forward neural networks, Chemometrics and intelligent laboratory systems 39, 43 (1997).
- Wang et al. [2020] Y. Wang, Q. Yao, J. T. Kwok, and L. M. Ni, Generalizing from a few examples: A survey on few-shot learning, ACM Comput. Surv. 53, 1 (2020).
- Natekin and Knoll [2013] A. Natekin and A. Knoll, Gradient boosting machines, a tutorial, Frontiers in neurorobotics 7, 21 (2013).
- Yang and Shami [2020] L. Yang and A. Shami, On hyperparameter optimization of machine learning algorithms: Theory and practice, Neurocomputing 415, 295 (2020).
- Dong et al. [2023] Q. Dong, L. Li, D. Dai, C. Zheng, Z. Wu, B. Chang, X. Sun, J. Xu, L. Li, and Z. Sui, A survey on in-context learning (2023), arXiv:2301.00234 .
- Gou et al. [2021] J. Gou, B. Yu, S. J. Maybank, and D. Tao, Knowledge distillation: A survey, International Journal of Computer Vision 129, 1789 (2021).
- Brown et al. [2020] T. B. Brown, B. Mann, N. Ryder, M. Subbiah, J. Kaplan, P. Dhariwal, A. Neelakantan, P. Shyam, G. Sastry, A. Askell, S. Agarwal, A. Herbert-Voss, G. Krueger, T. Henighan, R. Child, A. Ramesh, D. M. Ziegler, J. Wu, C. Winter, C. Hesse, M. Chen, E. Sigler, M. Litwin, S. Gray, B. Chess, J. Clark, C. Berner, S. McCandlish, A. Radford, I. Sutskever, and D. Amodei, Language models are few-shot learners (2020), arXiv:2005.14165 .
- Zhang et al. [2023b] Y. Zhang, Y. Li, L. Cui, D. Cai, L. Liu, T. Fu, X. Huang, E. Zhao, Y. Zhang, Y. Chen, L. Wang, A. T. Luu, W. Bi, F. Shi, and S. Shi, Siren’s song in the ai ocean: A survey on hallucination in large language models (2023b), arXiv:2309.01219 .
- Menard [2002] S. Menard, Applied logistic regression analysis, 106 (Sage, 2002).
- Han et al. [2021] X. Han, Z. Zhang, N. Ding, Y. Gu, X. Liu, Y. Huo, J. Qiu, Y. Yao, A. Zhang, L. Zhang, W. Han, M. Huang, Q. Jin, Y. Lan, Y. Liu, Z. Liu, Z. Lu, X. Qiu, R. Song, J. Tang, J.-R. Wen, J. Yuan, W. X. Zhao, and J. Zhu, Pre-trained models: Past, present and future (2021), arXiv:2106.07139 .
- Wang et al. [2023] Y. Wang, Y. Kordi, S. Mishra, A. Liu, N. A. Smith, D. Khashabi, and H. Hajishirzi, Self-instruct: Aligning language models with self-generated instructions (2023), arXiv:2212.10560 .
- Antonelli et al. [2019] L. Antonelli, M. R. Guarracino, L. Maddalena, and M. Sangiovanni, Integrating imaging and omics data: a review, Biomedical Signal Processing and Control 52, 264 (2019).
- Suthaharan [2016] S. Suthaharan, Support vector machine, in Machine Learning Models and Algorithms for Big Data Classification: Thinking with Examples for Effective Learning (Springer US, Boston, MA, 2016) pp. 207–235.
- Kudo and Richardson [2018] T. Kudo and J. Richardson, Sentencepiece: A simple and language independent subword tokenizer and detokenizer for neural text processing (2018).
- Pan and Yang [2009] S. J. Pan and Q. Yang, A survey on transfer learning, IEEE Transactions on knowledge and data engineering 22, 1345 (2009).
- Tal [2011] J. Tal, Chapter 14 - sample size, in Strategy and Statistics in Clinical Trials, edited by J. Tal (Academic Press, Boston, 2011) pp. 229–244.
- Liu et al. [2023b] P. Liu, W. Yuan, J. Fu, Z. Jiang, H. Hayashi, and G. Neubig, Pre-train, prompt, and predict: A systematic survey of prompting methods in natural language processing, ACM Computing Surveys 55, 1 (2023b).
- Yao et al. [2023] S. Yao, J. Zhao, D. Yu, N. Du, I. Shafran, K. Narasimhan, and Y. Cao, React: Synergizing reasoning and acting in language models (2023), arXiv:2210.03629 .
- Lewis et al. [2021] P. Lewis, E. Perez, A. Piktus, F. Petroni, V. Karpukhin, N. Goyal, H. Kttler, M. Lewis, W. tau Yih, T. Rocktaschel, S. Riedel, and D. Kiela, Retrieval-augmented generation for knowledge-intensive nlp tasks (2021), arXiv:2005.11401 .
- Vaswani et al. [2023] A. Vaswani, N. Shazeer, N. Parmar, J. Uszkoreit, L. Jones, A. N. Gomez, L. Kaiser, and I. Polosukhin, Attention is all you need (2023), arXiv:1706.03762 .
- Beltagy et al. [2020] I. Beltagy, M. E. Peters, and A. Cohan, Longformer: The long-document transformer (2020), arXiv:2004.05150 .
- cohere docs [2024a] cohere docs, Top-k and top-p, https://docs.cohere.com/docs/controlling-generation-with-top-k-top-p (accessed on January 29, 2024a).
- cohere docs [2024b] cohere docs, Temperature, https://docs.cohere.com/docs/temperature (accessed on January 29, 2024b).

### v2

@@ Discussion
We emphasize that classification cannot be successfully achieved using direct questioning alone (as seen in Supplementary Figs. S8-S10). Therefore, as part of our workflow, we designed a BERT-based classification model to categorize materials using text generated by the LLM. In practice, employing a classification model is a necessary step now (future, more materials robust, LLMs may not need this step), as directly querying the LLM yielded a near-zero accuracy rate.
The general workflow proposed here offers a novel approach to address a significant challenge in “AI for material science”, where datasets are often small. By utilizing a pre-trained BERT model (e.g., MatSciBERT optimized for material science), we can fine-tune the pre-trained model with textual data from a prompt-engineered LLM to achieve higher accuracy classification (up to 99.0% for the entire dataset).
We propose a general approach for materials classification and demonstrate its efficiency for a specific example, i.e., the classification of MGs. Our general workflow can be applied to any material applications, provided that labeled material samples are available. The findings point toward a future where integrating advanced language models and domain-specific fine-tuning will revolutionize material classification, especially in the very common cases where the data is sparse. We envision that LLMs may eventually serve as “world models” [49], encapsulating knowledge across diverse materials and domains. Consequently, such models based on text for classification, prediction, or even generating new materials will become potential tools for addressing challenges in material selection/design. A sufficiently robust LLM can provide desired classification outcomes directly through appropriate prompting and significantly streamline the process of materials discovery and design.
@@ Methods
We curated the original dataset for metallic glasses [40] by deleting of the entries (total 8,415) for which the original classification was ambiguous. The result was a dataset of 5,577 alloys and their respective classifications. To implement prompt engineering, we employed the langchain library [50] to systematically apply the template shown in Fig. 3 to each alloy in the dataset. We divided the 5,577 data points into a training and test set at a ratio of 80:20 to train the MgBERT model. The accuracy of the MgBERT model was assessed using the test set and compared with traditional machine learning models. We employed the Matminer [51] and Pymatgen [52] tools to convert the chemical formulas of alloys into “composition features” as defined within the Matminer [51] framework.
LLM and prompt engineering
@@
In this work, we employ the Google Gemini-pro [53] LLM, which surpasses human experts in massive multi-task language understanding [54]. As discussed earlier, our research harnesses the automated orchestration capabilities of the langchain [50] library to script the bulk generation of prompts and textual data. This scripting forms an integral part of a generic workflow and offers high adaptability. The prompt templates developed here are not rigid constructs but are designed to be inherently flexible, allowing for arbitrary modifications tailored to researcher demands and objectives. This level of customization is critical in materials science, where the nuances of the subject matter vary significantly from study to study. The ability to fine-tune prompts to align with particular research questions or datasets ensures that the language model can be effectively leveraged to generate practical and contextually relevant textual data outputs.
@@
Our investigation sought to demonstrate the effectiveness of prompt engineering by creating a range of alternative prompt templates to act as comparative benchmarks. These templates comprised one that integrated a chain-of-thoughts (CoT) approach with a few-shot learning paradigm (denoted “few-shot with CoT”), another that depended exclusively on few-shot learning examples (denoted “few-shot”), a third that was based on a zero-shot learning scenario (denoted “direct inquiry”), and our custom-designed “MetalPrompt” from our established workflow. By leveraging these diverse templates, we crafted specific prompts that were subsequently employed to interrogate the Gemini-pro LLM. The underlying rationale for this prompt strategy was to gather and juxtapose accuracy metrics for different prompting methodologies.
@@
Starting from MatSciBERT, we implemented a classifier function with MgBERT embedding to build MgBERT. Accuracy and cross-entropy are used to evaluate the model’s effectiveness. For classification within MgBERT, we used the value of the [CLS] token in the last layer of the attention section of MgBERT as the output to aggregate information for the entire model. Then, through a classifier, we compress the value of the token into a vector of length 3, and label 0 as BMG, 1 as Ribbon, and 2 as Non-ribbon. We then identified the maximum position of the output vector value and used this as the final classification result. The overall expenses associated with MgBERT training was $46 USD on an Nvidia V100 (i.e., 30 hours at $1.53 USD/hour). The cost of 1.89 million input and 2.14 million out tokens was $1.5 USD (official Gemini-pro pricing on January 29, 2024). To apply MgBERT for inference on 1,000 different compositions cost $0.3 USD. This training and application of MgBERT enables researchers to cost-effectively apply this general approach across a wide range of materials applications.
@@
To elucidate input and output relationships, we applied the LIME [44] (local interpretable model agnostic explanations) technique; widely used to explain predictions of machine learning models by constructing relationship functions via interpolation sampling. Here, we displayed the 25 most important features and performed 1,000 interpolation samples for function fitting.
To analyze the input-output relationship at the attention level, we used the bertviz [55] library to complete the attention score flow diagram for 12 layers and 12 attention heads. To attribute the importance of the [CLS] token in the final attention layer, we first averaged the attention score of the 12 attention heads in the last layer, then divided the input text words into three layers based on their positions (i.e., composition, element, and alloy). To evaluate the importance of each layer, we calculated the sum of attention scores within that layer, weighted by the token length ratio, and then divided by the total token length.
@@
The method section provides the models and algorithms employed in this study, while specific parameter implementations can be found in the Supplementary Information (SI).
@@
All the codes used in the paper will be made available on GitHub upon the acceptance of manuscript, and these codes will be provided upon reasonable requests.